Solve by the substitution method
9x-4y=-85
8x+96=y

Answer by checkley77(12844)

9x-4y=-85
8x+96=y
9x-4(8x+96)=-85
9x-32x-384=-85
-23x=-85+384
-23x=299
x=299/-23
x=-13 ans.
9*-13-4y=-85
-117-4y=-85
-4y=-85+117
-4y=32
y=32/-4
y=-8
8*-13+96=-8
-104+96=-8
-8=-8